3. For hypo-eutectoid steel (say, carbon content equal to 0 .40 wt%), if cool from austenite to below eutectoid temperature, calculate the relative amount of pro-eutectoid ferrite and Pearlite. What is the carbon content of the Pearlite microconstituent? What is the carbon content of the pro-eutectoid constituent? Calculate the relative amount of a and Fe3C phase in the perlite constituent. 

4. For hyper-eutectoid steel (say, carbon content equal to 1.2 wt%), if cool from austenite to below eutectoid temperature, calculate the amount of pro-eutectoid Cementite and pearlite. What is the carbon content of the proectectoid cementite constituent? Calculate the relative weight fraction of α and Fe3C in the sample.
Answer #1

3. For hypoeutectoid steel of composition 0.40% wt C.

Amount of pro-eutectoid ferrite is = (0.8 - 0.40)/(0.8 - 0.002)

= 0.40/0.798)

= 0.5012 or 50.12%

In pearlite, carbon % is 0.83 wt%C

In pro-eutectoid steel carbon % is greater than 0.80 wt%C

In pearlite

Relative amount of alpha iron is (6.67 - 0.83)/(6.67 - 0.002)

= 0.8758 or 87.58%

And relative amount of cementite is = 1 - 0.8758 = 0.1241 or 12.41%.

4. Amount of pro-eutectoid cementite in 1.2 wt% C hyper eutectoid steel is = (1.2 - 0.83)/(6.67 - 0.83)

= 0.0633 or 6.33%

Amount of pearlite in hyper-eutectoid steel is

= 1 - 0.0633 = 0.9267 or 92.67%

Carbon content of pro-eutectoid cementite is 6.33%.

  

=